For the following reaction, 14.4 grams of phosphorus (P4) are allowed to react with 36.5 grams of chlorine gas

phosphorus (P4)(s) + chlorine(g) phosphorus trichloride(1)

What is the maximum mass of phosphorus trichloride that can be formed? grams

What is the FORMULA for the limiting reagent?

What mass of the excess reagent remains after the reaction is complete?

grams

Submit Answer

Retry Entire Group

2 more group attempts remaining
1
Expert's answer
2021-10-10T10:28:11-0400

P4+6Cl2→4PCl3

Molar mass of phosphorus = 30.973762

14.4/30.973762= 0.465

= 0.465× 4= 1.86moles

Molar mass of Cl2 = 70.9

36.5/70.9= 0.51481mol

= 0.51482× 4/6

= 0.343mol

= 1.86+0.343= 2.203

Molar mass of PCl3 = 137.33 g/mol

137.33×2.203

= 302.57grams